### v1

# Enhancing multimodal analogical reasoning with Logic Augmented Generation

Abstract
Recent advances in Large Language Models have demonstrated their capabilities across a variety of tasks. However, automatically extracting implicit knowledge from natural language remains a significant challenge, as machines lack active experience with the physical world. Given this scenario, semantic knowledge graphs can serve as conceptual spaces that guide the automated text generation reasoning process to achieve more efficient and explainable results. In this paper, we apply a logic-augmented generation (LAG) framework that leverages the explicit representation of a text through a semantic knowledge graph and applies it in combination with prompt heuristics to elicit implicit analogical connections. This method generates extended knowledge graph triples representing implicit meaning, enabling systems to reason on unlabeled multimodal data regardless of the domain. We validate our work through three metaphor detection and understanding tasks across four datasets, as they require deep analogical reasoning capabilities. The results show that this integrated approach surpasses current baselines, performs better than humans in understanding visual metaphors, and enables more explainable reasoning processes, though still has inherent limitations in metaphor understanding, especially for domain-specific metaphors. Furthermore, we propose a thorough error analysis, discussing issues with metaphorical annotations and current evaluation methods.
Index Terms:
Logic Augmented Generation,Figurative Language Understanding,Analogical Reasoning,Large Language Models,Conceptual Blending Theory
I Introduction
From musical improvisation to solving complex scientific problems, much of what we know and do relies on implicit knowledge: the vast, unconscious reservoir of information we acquire through experience rather than formal instruction [^1] Implicit and tacit knowledge are usually treated as synonyms, for a more comprehensive account on the distinction between the two see [1].. This type of knowledge is employed in analogical reasoning, when we rely on stored patterns, schemas, and relational structures to retrieve and apply knowledge that we may not have explicitly articulated. Therefore, implicit knowledge shapes how we interpret the world, draw inferences, identify conceptual links, and even understand figurative language, all without deliberate effort [2]. Unlike humans, however, machines struggle with these tasks, as they lack direct experience of the physical world, making the computational modeling of implicit learning and consequently analogical reasoning a significant challenge [^2] As shown by the emergence of dedicated conference workshops on the matter, such as the Workshop on Analogical Abstraction in Cognition, Perception, and Language (Analogy-Angle): https://analogy-angle.github.io/.
Figure 1: The LAG framework. The input can be a text or an image, then converted into a FRED-like KG. The graph is then extended with LLM-generated triples according to specific heuristics and defined tasks.
Recent work considers metaphor the highest level in the analogical taxonomy and presents the greatest challenge for Large Language Models (LLMs) [3]. This difficulty arises from the absence of effective prompting techniques to capture the inherent compositionality of language, as well as the lack of anchoring to commonsense knowledge and visual information [4]. Instead of surface-level similarities, analogical reasoning requires the understanding of relational links between concepts, a skill usually learned through the physical interaction with the world [5, 6]. The lack of understanding of relational links in LLMs is a reasoning deficit that remains undetected by standard benchmarks [7], particularly with respect to metaphors. This issue also arises from the lack of a shared problem definition for metaphors and a scarcity of robust, heterogeneous data and evaluation strategies [8].
Although LLMs are trained on diverse corpora that may contain fragments of implicit meaning embedded in human language, analogical reasoning on implicit knowledge is not explicitly codified during training, but might still be represented within LLMs as probabilistic associations across continuous vector spaces. However, relying solely on the probability of association of words to understand and generate language restricts current LLMs [9].
Therefore, despite the multiple remarks that AI systems should be enriched to better mirror deeper semantic and cognitive models of the world, rather than merely relying on correlations between words [10, 11, 12, 13], analogical reasoning remains a significant struggle for the field of Natural Language Processing [14, 13].
Conceptual Blending Theory (CBT) models meaning construction through partially compositional processes in analogy, metaphor, counterfactuals, and related phenomena [15, 16]. Despite its computational promise, studies on CBT largely overlook metaphorical aspects [17, 18, 19] because systematically encoding the vast, unbounded encyclopedic knowledge required for metaphors is challenging [8]. Furthermore, the absence of a core blending ontology has so far limited the computational understanding of these cognitive processes that are, for the most part, innate and implied. Indeed, research shows that LLMs’ performance improves significantly with structured background material [20, 21]. Consequently, combining reliable but labor-intensive semantic technologies with faster, context-dependent yet error-prone LLMs—as demonstrated by Logic Augmented Generation (LAG) in the work by Gangemi and Nuzzolese [22]—opens a promising new avenue for metaphor understanding and analogical reasoning.
This paper proposes a framework based on LAG and inspired by CBT to extract extended Knowledge Graph (KG) triples containing implicit meaning from unlabelled metaphorical data in text and image format (Fig. 1). Integrating structured, machine-interpretable semantics through KGs with fast and dynamic LLM systems allows the real-time flexible understanding and representation of the implied aspects of metaphor in data across various domains and modalities such as news, visual advertisements, memes, and medical sciences.
The contributions of this work are: (i) the adaptation of the LAG framework to enhance multimodal analogical reasoning; (ii) its evaluation on variations on the task of metaphor detection and understanding, namely: linguistic metaphor detection, conceptual metaphor detection and understanding, and visual metaphor understanding; and (iii) a discussion on its computational implementation and applications [^3] All supplementary material, code and data are available at https://github.com/dersuchendee/knowledge-augmented-generation.
The structure of the paper is as follows: Section I introduces the significance and contributions of this work; Section II provides an overview of the theoretical frameworks of analogical abstraction in metaphors; in Section III, we introduce the Blending Ontology for metaphorical conceptual blending. Section IV describes the LAG methodology employed in the paper, Section V and VI concerns its application to three tasks. Section VII analyzes the results, and Section VIII summarizes the contributions and implications for extracting implicit meaning to reason analogically from figurative unlabelled data.
II Background
In this section, we explore the background of our model and discuss the challenges related to existing automated processes for analogical abstraction in the context of metaphor.
II-A Implicit knowledge and analogical abstraction
Implicit learning is a process by which knowledge is acquired independently of conscious attempts to do so [23].
Analogical abstraction is a kind of implicitly learned knowledge where as the learner detects and reproduces patterns or analogies across experiences or concepts. This recognition of structural similarities in the process of analogical reasoning can be used to solve problems and make accurate decisions about novel circumstances, and is particularly clear when the connections between concepts are not obvious or are learned through repeated exposure.
A fundamental component of NLP for many years, LLM-based analogical reasoning has been recently investigated. Webb et al. show that LLMs excel at identifying simple proportional word analogies [24]. However, other studies show LLMs’ performance diminishes when faced with more complex rhetorical figures [3], such as metaphors, or when the task requires distinguishing between associative and relational responses [25]. These findings suggest that while LLMs can identify associations learned from data, they do not necessarily grasp the underlying relationships that define how concepts are connected in more abstract ways. As a result, there is a need for a structured framework that directs these systems to undertake specific analytical steps during tasks that require analogical abstraction while making space for explainability.
II-B Conceptual Blending Theory, Metaphor, and Frame Semantics
According to Conceptual Metaphor Theory [26], metaphors create parallels between domains, highlighting experiential correlations. This claim, along with with Mental Spaces Theory, is expanded by Conceptual Blending Theory (CBT), which describes one of the basic cognitive operations of human beings: conceptual combination. According to CBT, in metaphor, a generic space makes abstract and links the two source and target input spaces and generates a new blended space according to a metaphorical criterion or “property” [15]. Indeed, many studies [27, 28, 29] claim that by explaining one object in terms of another, the ‘target’ domain receives a property from the other object, the ‘source’, or in general that the source and target are associated in thought and language because they share a property of some kind. For example, in the metaphor ideas are food, ideas and food are connected through the criterion of internalization (food is inserted and processed in our bodies like ideas), and the two spaces are abstracted with respect to the agent, the input, the process, the output, selectively basing on the context. In this way, the connecting property (internalization) is a higher-level concept that transcends the inherent characteristics of each input space and motivates the comparison.
However, the existence of a shared property is by no means a sufficient condition for two concepts to participate in a metaphorical mapping [30]: for metaphorical meaning to bridge between individuals and within a community, those individuals must already share some understanding of the components involved in the metaphor [31] due to commonsense knowledge and/or specific cultural or social context in which a metaphor is used [32]. In this regard, the CBT framework can be explicitly connected to Fillmore’s frame semantics [33], where certain verbs automatically invoke particular roles, values, and perspectives of discourse participants [34]. Here, different communicative contexts activate background knowledge (i.e. frames) that constrains interpretation. For instance, as shown in Fig. 2, when the conceptual metaphor ideas are food is invoked in the sentence “This is what happens when science serves up facts more indigestible than evolution”, people must already have a background to draw from, some individual grasp of both scientific knowledge and food to understand that it means a way science proposes an idea that is hard to accept from the speaker’s point of view, and that it is a negative outlook on the two domains. The focus-refocus ability of the metaphor is therefore strongly connected to these theoretical frameworks.
Figure 2: The metaphor ideas are food as shown in an example sentence present in the Framester knowledge base [35]. Here, not all roles that exist for each frame mutually map (those that map are in bold and connected through a dotted line), only some are activated in the utterance. Active roles are then subsumed by the roles of the Blending frame (the property), internalization. Overall, the information about the frames is part of the background of the utterance, while the attitude through which we see it is negative.
II-C Theory-integrated metaphor detection and understanding
Recent advancements in metaphor detection have increasingly incorporated theoretical frameworks into computational approaches. Studies leveraging Conceptual Metaphor Theory (CMT) have shown improved baseline results in metaphor identification [36, 37]. For metaphor interpretation, various methods have been developed to discern the source domain implied by the metaphor, given a target domain. These include unsupervised techniques [38], deep learning-based approaches [39], and methods using Large Language Models [40]. However, these methods generally depend on annotations of a single domain rather than exploiting unlabeled data. In Hicke et al., [8], when unlabeled data was exploited, the Metaphor Identification Procedure was preferred as a theoretical framework, thus the metaphors were annotated by category but not by source and target domain. Indeed, the authors recognize this limitation, stating that with the Metaphor Identification Procedure, forcing the models to annotate word-by-word makes it challenging for them to identify metaphors comprising multi-word units.
Other research focuses on extracting properties linking the source and target domains to interpret metaphors [41, 42, 43]. While these approaches yield promising results, they often do not adequately capture the underlying theoretical reasoning processes that are based on the definition of their structural similarity in terms of abstraction and relational dependency on the two metaphorical domains. Additionally, for these systems to work, they require a specified target word.
In the context of visual metaphor understanding, while the work by He et al. [44] focuses on detection, it does not adequately address metaphor understanding according to existing theoretical frameworks, according to which a comprehensive approach should at least involve identifying both source and target domains. To foster richer research in visual metaphors, the Metaclue task set was introduced [45], though its code and dataset are not publicly available. Metaphoric compositions linked by an annotated attribute are available in the ELCo Dataset proposed by Yang et al. [4], where metaphorical composition in emojis is evaluated via a textual entailment task. Integrating heuristics based on existing ontologies on metaphor and knowledge graphs into the metaphor detection and understanding pipeline remains largely unexplored, despite its potential to render theories standardized and machine-readable across multimodal contexts, reduce hallucinations, adapt them for other automated reasoning tasks, and enhance system explainability.
II-D Logic Augmented Generation
The so far described limitations of theory-integrated metaphor detection and understanding studies can be addressed with the integration of ontologies and commonsense KGs in an LLM-based metaphor detection pipeline. The Logic Augmented Generation (LAG) approach, as shown in Gangemi and Nuzzolese [22], conceptualises LLMs as potential Reactive Continuous Knowledge Graphs (RCKGs), which can dynamically adapt to diverse inputs by extending and contextualising Semantic Knowledge graphs (SKGs) that work as base models. LAG exploits the duality between implicit knowledge generated during training and explicit knowledge, externalised in the form of structured knowledge graphs through generation and prompting. RCKG extraction involves a three-step process, i.e. (i) mapping multimodal signal to natural language, (ii) converting natural language into an SKG, and (iii) extending the SKG with tacit knowledge based on multiple heuristics (the extended knowledge graph as XKG).
In this frame of reference, SKGs ensure logical consistency, enforce factual boundaries, and foster interoperability, while LLMs process unstructured data and provide contextual insights on implicit knowledge on demand.
III The Conceptual Blending Ontology
The Conceptual Blending ontology [^4] Available at http://www.ontologydesignpatterns.org/ont/mdns/blending.owl serves as a heuristic frame of reference in the LAG process. It is based upon mDnS (mereological Descriptions and Situations) [^5] http://www.ontologydesignpatterns.org/ont/mdns/mdns.owl, the Cognitive Perspectivisation ontology as defined in Gangemi and Presutti  [46] [^6] http://www.ontologydesignpatterns.org/ont/persp/perspectivisation.owl, and the MetaNet schema from the Amnestic Forgery ontology by Gangemi et al.  [47].
III-A Background
mDnS is a novel axiomatization for the D&S ontology pattern [^7] http://www.ontologydesignpatterns.org/cp/owl/descriptionandsituation.owl, fully compatible with classic D&S and constructive D&S (cDnS) and is based on intensional/extensional mereology and optimizing expressivity with respect to low complexity for a hybrid reasoner. By introducing the concepts of descriptions and situations, the Description and Situations ontology design pattern  [48] provides a formalization of frame semantics where a situation is a set of facts as described by an observer that is interpreted by a description, a schema that defines the entities observed in a situation.
Situations can be reported from different perspectives: the Cognitive Perspectivisation Ontology  [46] represents the Perspectivisation frame, a type of situations where a fact is reported within a certain narrative, or Lens, which creates a viewpoint, cutting the reality which a Conceptualiser holds an attitude towards. Perspectivisation itself is therefore not merely rhetorical, as it blends entities playing two roles: the cut and the lens. Finally, the Amnestic Forgery ontology  [47] is a model for metaphor semantics based on MetaNet and compatible with Conceptual Metaphor Theory [49]. Amnestic Forgery reuses and extends the Framester schema, allowing to deal with both semiotic and referential aspects of frames.
III-B Ontology
The proposed ontology generalises the notion of metaphors as frame mappings, and uses mDnS for ground axiomatisations. The ontology defines key classes and properties that operationalize the blending of concepts:
  • Classes such as Blend, Blendable, Blended, and Blending represent different stages and elements of the conceptual blending process.
  • Object Properties such as blendableComponent,
    blendedComponent, and blendingComponent are used to link these classes in meaningful ways that reflect the underlying cognitive processes. These properties help in mapping the components of one conceptual domain to another, leading to the emergence of new, blended concepts. Other properties like enablesBlending and inheritsRoleFrom define the interaction between different conceptual elements within the ontology.
More specifically, as shown in Fig. 3, the basic Descriptions include:
  1. 1.
    Blending, a generic description enabling Blendable frames to be mapped by subsuming some of their components;
  2. 2.
    Blendable, two descriptions with components that are subsumed by components in the Blending frame;
  3. 3.
    Blended, the description that inherits components from the Blendable frames, firstly through the Blending sharable space, and eventually inheriting other components from the Blendable frames, giving rise to a novel set of blended situations;
  4. 4.
    Blend, the meta-description that has as components Blending, Blendable, and Blended.
Figure 3: The core of the Blending ontology [^9] The full diagram can be seen on Github https://github.com/dersuchendee/knowledge-augmented-generation.
When a conceptual metaphor is formulated from two input frames, their roles are generalized and inherited (a Blendable inheritsRole) from a more generic Blending, which enables the blending (EnablesBlending). The Blended in turn inherits components from the Blendable frames through the Blending generic space. The Blended is therefore a new, independent space where the metaphor takes place.
IV Methodology
The proposed method, based on LAG [22], aims to enhance the detection and understanding of metaphor by establishing a standardized and explainable model for the semantic representation of implicit natural language. Therefore, our approach has two main components: i) the representation of inputs, regardless of their modality, into natural language, ii) their conversion into a base knowledge graph with Text2AMR2FRED and iii) the SKG enhancement driven by Conceptual Blending Theories applied to metaphor.
IV-A Automated natural language and image KG representation with Text2AMR2FRED
Text2AMR2FRED has been first introduced as a revised architecture of FRED’s text-to-KG pipeline  [50], using state-of-the-art Abstract Meaning Representation (AMR) parser with multilingual capabilities. In this way, relevant commonsense knowledge from knowledge bases linked by the Framester knowledge base [35] is retrieved and represented in a graph.
Figure 4: Text2AMR2FRED pipeline for the sentence “Crime has infected communities everywhere”.
For example, as shown in Fig. 4, the sentence “Crime has infected communities everywhere” is first parsed into an AMR graph, using SPRING  [51]. The AMR graph is then converted into an RDF/OWL knowledge graph that follows FRED’s knowledge representation patterns using AMR2FRED.
AMR identifies the cause of the infection in the sentence, which is, however, not a germ, but crime and the thing that gets infected (community) everywhere. In the FRED-like KG, this graph is enriched with knowledge from Framester such as frames, WordNet synsets and PropBank roles, along with alignment to the DOLCE foundational ontology  [52]. However, there is no representation of metaphorical knowledge yet at this stage. While a conceptual metaphor KG sourced from Metanet is contained in Framester, it is still necessary to be able to detect figurative language in any text. An image can also be used by being automatically described in natural language. In this case, language is an amodal way to bridge modalities in the resulting supramodal SKG. The SKG can then be extended with implicit contextual knowledge with the support of an LLM.
IV-B Logic Augmented Generation with heuristics
Figure 5: Logic augmented graph for the sentence “Crime has infected communities everywhere”. In black is the base graph, in red the augmented classes and properties.
As shown in Fig. 5, the result of the framework that generates XKG triples is the merging of the original KG with new classes and properties according to a specific ontology in the prompt. In this case, a minimum requirement of selected classes of the Blending Ontology is specified to account for the combination of concepts. Considering this framework, it is possible to use Conceptual Blending specifically for metaphor, where the two domains, the source and the target, correspond to the Blendable frames. These frames have an overarching Blending frame in common that both explains and motivates the metaphor. Such fine-grained element mappings are required to understand the consistency of the blend, something that has been mentioned in research but so far has yet to be implemented [40]. The Blending frame motivating the metaphorical blend also makes a system’s analogical reasoning capabilities more explainable. For example, a system can correctly identify the elements involved in a metaphor but misidentify the property that links them, as discussed in Section VII, showing the inability to understand the metaphor. With this framework in mind, it is possible to create a prompt for an LLM that, given a text and the resulting KG representation from Text2AMR2FRED, it extends it to create an augmented RDF graph in Turtle format.
V Application and Evaluation
In this section, we evaluate the framework’s ability to extract implicit meaning by applying it to metaphor detection and understanding. First, we test the LAG framework on three metaphor detection datasets. The best configuration is then applied to three metaphor understanding datasets. Finally, we perform ablation tests on metaphor detection and visual metaphor understanding.
V-A Tasks
This section shows the application of the framework in three metaphor-related tasks: i) metaphor detection, where a SKG derived from a sentence in natural language is enhanced to determine if a sentence is metaphorical, sometimes with a target word as a reference; ii) conceptual metaphor understanding, which additionally aims to identify the correct source and target domains that make up the metaphor, and iii) visual metaphor understanding, where the aim is to identify the property linking the two domains in the visual metaphor.
V-A1 Metaphor detection and understanding
As shown in the Supplementary Material, the LLM is given a sentence, a SKG representing the sentence derived from Text2AMR2FRED, and a series of instructions. The prompt first requires identifying the possible source and target of the sentence; then, it is asked to define these two as conceptual frames containing roles and following a blending function that orchestrates the creation of a blended space. Following the Blending Ontology, a Lens and an Attitude further refine the definition of the metaphor. An example of the graph for a specific sentence is also provided.
Given the Text2AMR2FRED graph, and the presumed source and target of the sentence given a target word when available, along with blending-derived instructions, the LLM is prompted to analyze the sentence and to use at least these specific ontology elements according to the given definitions, to determine the possible metaphoricity of a sentence:
  1. 1.
    bl:Blending,
  2. 2.
    bl:Blendable,
  3. 3.
    bl:Blended,
  4. 4.
    cp:Attitude,
  5. 5.
    cp:Lens,
  6. 6.
    the datatype property metanet:isMetaphorical, which classification is based on.
Furthermore, in the prompt it is asked to determine blendable roles and their mapping to create a blended metaphorical space according to the Blending Ontology.
V-A2 Visual metaphor understanding
Following the original study performed on human participants outlined in Petridis et al.  [32], the methodology for visual metaphor understanding involves the presentation of three examples and correct annotations of visual metaphor elements within an in-context learning technique, along with a new image to analyze. In this case, the image to analyze is first described in natural language, and the text is passed to Text2AMR2FRED to generate a base graph. The base graph, image, accompanying sentence, and blending instructions are all included in the prompt to generate the augmented KG, as shown in Fig. 6. In this work, we experiment with multiple combinations of these inputs to find the best prompting technique.
Figure 6: Example prompts and extracts of correct and wrong outputs according to manual evaluation for visual metaphor understanding.
VI Datasets
For the purpose of this study, and for comparison across the outlined tasks, for textual metaphor detection we conducted experiments on two publicly available metaphor datasets: (1) MOH-X [53], which comprises 647 sentences annotated as either metaphorical or literal; and (2) TroFi [54], another metaphor detection dataset collected from the 1987-1989 Wall Street Journal Corpus. We sampled 300 balanced instances each for testing. For conceptual metaphors, we employ (3) the dataset refined by Wachowiak and Gromann [40] of 447 sentences including conceptual metaphors annotated with source and target domain (for ease of reference, we will call it WG). We also introduced a fourth dataset to verify the capability to identify and explain domain-specific conceptual metaphors: (4) the Balanced Conceptual Metaphor Testing Dataset (BCMTD). This dataset consists of 147 sentences, equally distributed among general conceptual metaphors (derived from annotated Metanet-derived data from Framester [35] and that weren’t present in WG), scientific conceptual metaphors (extracted from scientific literature from the Appendix in Van Rijn and Van Tongeren [55] and examples from Semino et al. [56], allowing a domain-specific dataset on medical metaphors), and literal sentences from the VUA corpus [57] (previously refined by Wachowiak and Gromann [40] as many of the sentences were wrongly labeled or deemed unfit for the task). The visual metaphors dataset is derived from Petridis et al. [32]. It contains 48 images, 27 ads and 21 non-ads, and three examples, that have been given to human participants to test their ability to identify visual metaphors according to a conceptual blending framework. Table I shows statistics for all the datasets.
TABLE I: Summary of datasets statistics, including instances, percentage of metaphorical sentences, and samples.
Dataset # Instances % Met. # Samples
MOH-X 647 48.7 300
TroFi 3737 43.5 300
WG 447 100 447
BCMTD 170 (conceptual) 451 (scientific) 50 (VUA) 66.6 147
Visual metaphors 51 100 48
VI-A Evaluation metrics
The evaluation metrics for metaphor detection are both automated and manual depending on the type of task.
VI-A1 Automated evaluation metrics
We use Accuracy and F1 score to evaluate the model’s performance both in generic and conceptual metaphor detection. For conceptual metaphor understanding, evaluated via the correct identification of both source and target mappings according to a gold standard, the BLEURT score [58] has been shown to capture semantic similarities between texts. It is considered in order to evaluate the distance between the candidate and reference source and target in the metaphor. This is an evaluation method similar to the one in Saakyan et al. [59], where instead of only reporting label accuracy, it is reported according to defined thresholds. In this case, the threshold is where the correctly predicted labels of source and target domains achieve a BLEURT positive score.
VI-A2 Manual evaluation
To correctly determine metaphorical source and target mappings for domain identification, and to correctly compare gold standard and predicted blending properties for visual metaphor interpretation, it is also necessary to conduct a manual evaluation. The best performing solution for conceptual and visual metaphor understanding is therefore evaluated for source and target mappings by three annotators. The annotators were graduated in linguistics, had proven expertise on metaphors, and had English Native or near-native language proficiency. They exhibited a Fair inter-annotator agreement (Fleiss’ Kappa) in all cases, with an average on all tasks of 0.35, which is understandable for a task so subjective as metaphor understanding. More information on the agreement scores is in the Supplementary Materials. To calculate the Accuracy, we compute an average of each Accuracy score. In the visual metaphor interpretation task, the manual evaluation compares the gold standard property with the predicted one connecting source and target. The evaluation directly follows the one in Petridis et al. [32], which considered as correct responses directly related to the objects in the image. We also repeated the study on a sample of 24 images on 20 participants of the University of Bologna, manually evaluating their responses as well.
VI-B Baseline methods
In Tian et al. [37], several baselines for metaphor detection are discussed, with the best-performing method being a Theory Guided Scaffolding Instruction Framework based on Conceptual Metaphor Theory (TSI CMT). MetaPRO is a baseline among non-LLM methods for detecting source and target domains, serving as the state-of-the-art approach for identifying mappings without specifying a target word [36]. For conceptual metaphor understanding, no current baselines simultaneously detect source and target but only one of the two. We thus employ LLMs in different configurations, including zero-shot and few-shot, shown to be effective in previous research [40]. No established baselines for visual metaphor understanding specify source, target, and blending properties. Thus, we explore different configurations of the LAG framework using the same three few-shot examples used to test humans in [32], and evaluate against the few-shot baseline without LAG.
VI-C Parameters setup
In our experiments, Claude 3.5 Sonnet [^10] https://www.anthropic.com/news/claude-3-5-sonnet. serves as the primary LLM. We compare the best-performing results with Llama 3.1 70B Instruct Turbo [^11] https://ai.meta.com/blog/meta-llama-3-1/ in all cases except for visual metaphor interpretation, where no current open models support few-shot image examples. Only the APIs are used for testing. We conducted tests using default parameters and set the temperature to 0 for greedy decoding.
VI-D Main results
In this section, we outline the main results of the LAG framework applied to the selected tasks.
VI-D1 Metaphor detection: comparison with baselines
Table II shows comparative results of LAG compared with baselines for the sampled TroFi and MOH datasets. Table III shows comparative results of LAG with respect to the BCMTD.
TABLE II: Performance comparison of various methods for metaphor detection on MOH-X and TroFi. Best performing results are in bold.
Method MOH-X TroFi
F1 (%) Acc. (%) F1 (%) Acc.(%)
MetaPRO 84 81 79 70
TSI CMT* 82.5 82.9 66 66.8
LAG 89.7 87.3 89.7 84.6
TABLE III: Performance metrics for the BCTMD Dataset for metaphor detection. Best performing results are in bold.
Method Accuracy (%) F1 Score (%)
LAG 80.1 84.1
MetaPRO 69.1 69.8
Few-Shot 12 59.0 48.9
Few-Shot 6 52.4 45.2
Few-Shot 3 47.5 42.8
Zero-shot 22.9 33.8
VI-D2 Conceptual metaphor understanding
Conceptual metaphor understanding is a more challenging task than simple metaphor detection, as it involves the correct detection of both source and target domains for a metaphor, which shows whether the system has understood it or not. We employ majority voting between the three annotators to select the gold standard label.
WG dataset. Manual evaluation performed by the three annotators for the dataset reveals that the percentage of metaphors where source and target were recorded to be correctly identified at the same time is 25.6%. BLEURT scores showed a moderate correlation between the human annotations and the predicted domains. For the target domain, a point-biserial correlation of 0.702 (p <<< 0.05) and a Spearman correlation of 0.698 indicate that BLEURT is reasonably sensitive to the human judgments. For the source domain, a point-biserial correlation of 0.707 (p <<< 0.05) and a Spearman correlation of 0.680 were observed.
BCTMD. Manual evaluation performed by the three annotators for the best performing methodology in metaphor understanding, in fair agreement, reveals that out of 96 conceptual metaphors, the percentage of correctly identified source and target at the same time is 34.3%. Furthermore, the distribution of correctly identified metaphors is 51.6% for generic conceptual metaphors, while for scientific conceptual metaphors the score lowers to 8%. This score, with respect to the generic metaphors also contained in WG, might be related to the fact that there is no domain-specific metaphor dataset on which LLMs have been trained, but also that scientific metaphors have different levels of conventionality, tending to result more literal as they come to be widespread in language. BLEURT scores showed a 36.9% positive correlation between reference and predicted source and target domains. A point-biserial correlation of 0.3525 (p <<< 0.05) and Spearman correlation of 0.2917 (p <<< 0.05) both indicate moderate positive correlations for BLEURT.
VI-D3 Visual metaphor understanding
For visual metaphor understanding, the identification of the blending property corresponds to the correct interpretation of the source and target employed in the metaphor (as shown by Petridis et al.[32]). Out of the three configurations, shown in Table V, manual annotation from the three evaluators revealed that the best performing result is LAG without the insertion of a sentence describing the image alongside the image itself in the prompt. Concerning human evaluation described in Petridis et al. [32], which correctly interpreted the meaning of the image 41.32% of the time, in the best performing method the LLM correctly interpreted the meaning an average of 67.06% of the time. We then sampled the dataset, randomly taking 24 visual metaphors out of the 48 in the dataset, and repeated the original study by [32] with 20 human participants recruited from the University of Bologna, from student to researcher level, with native or near-native English fluency. This setup slightly differs from the original study, where participants were employed without specifying their study level. Participants’ submissions were anynomous and informed consent was provided. This time as well, on average, human participants have correctly interpreted the meaning 59.25% of the time, with respect to the average LLM score of 73.5% for the same images.
TABLE IV: Comparison of methods for visual metaphor understanding. The best performing result is LAG without the injection of a sentence that describes the image in the input.
Method Avg. Accuracy (%)
LAG sent+img 65
LAG no sent 67
LAG no img 65.2
Few-Shot (3) 54.7
TABLE V: Ablation study results for the visual metaphor dataset. Best results in bold.
Method Avg. Acc.
LAG no sent 67
No Blending 68.6
No Graph 56.2
VI-D4 Metaphor detection and understanding: comparison with open-source LLM
We compare the best performing results of metaphor detection with Llama, 3.1 70B Instruct.
TABLE VI: Comparison of LAG for metaphor detection in each dataset with Claude and with Llama. Best results in bold.
LLM MOH-X TroFi BCMTD
Acc. F1 Acc. F1 Acc. F1
Claude 87.3 89.7 84.6 89.7 80.1 84.1
Llama 55.9 69 60.8 75 66.6 66.6
For what concerns conceptual metaphor understanding, according to the evaluators, the average percentage of Accuracy is 37.5% against the method with Claude, which yields an Accuracy of 34.3%. The results show a drop in performance when using Llama with metaphor detection, but metaphor understanding of both source and target mappings seems to perform slightly better with respect to Claude.
VI-D5 Ablation studies
We conduct the ablation study to evaluate the impact of the components in our method for metaphor detection by two configurations: (i) removing the blending instructions, and ii) removing the knowledge graph injection in the prompts. All the prompts used in these configurations are in the Supplementary Materials. Table VII shows the results of the ablation studies for metaphor detection for all the textual datasets, while Table V for visual metaphor understanding.
TABLE VII: Ablation study results for MOH-X, TroFi, and BCMTD datasets. Best results in bold.
Method MOH-X TroFi BCMTD
Acc. F1 Acc. F1 Acc. F1
LAG 87.3 89.7 84.6 89.7 80.1 84.1
No Blending 81.6 87 81.9 86 78.6 85.2
No Graph 78.6 82 83.9 87 70 73
VII Discussion
In this section, we discuss the results, outline some applications and further research directions for the proposed framework, and investigate limitations.
VII-A Application of LAG to multimodal metaphor detection and interpretation
The results outlined in Section V show that for the task of metaphor detection, the augmented KG according to a prompt and a base graph that constrain the LLM to adhere to the Blending Ontology outperforms the current metaphor detection and understanding baselines. This approach can be used to extract knowledge graphs from unstructured multimodal metaphorical data. As such, it could have effects in different applications, for example, improving hate speech, which often has metaphorical features [60], and analyzing debates in the press and online media. For what concerns visual metaphors, since humans tend to struggle to identify them, the automated understanding of visual blends, for example in advertisement, can bridge this gap. Additionally, using an open-source LLM like Llama can decrease the performance concerning detection, but not understanding tasks. Ablation studies show that while removing the graph reduces the performance in all cases, removing the blending heuristics in the prompting reduces it in all cases besides visual metaphor understanding. While this is a slight difference, it could be motivated by the fact visual metaphors can be wrongly described in text before processing. Another explanation for this slight performance drop is shown by the error analysis (see Section VII-B), where incorrect property was the most frequent issue both for humans [32] and LLMs: the objects are detected in the images, but there is still a struggle with extracting the correct properties. Because half of the images are advertisements, which typically present clearer context, the intended properties of the objects are usually clearly defined. By not specifying the purpose of the image and its context, we make it challenging for both humans and machines to accurately determine the metaphorical property via a blending mechanism that tends to be more open-ended without contextual specifications.
VII-B Quality of the augmented graphs and error analysis
We provide a sample of 16 generated graphs, 8 derived from correctly identified metaphorical sentences, 8 derived from visual metaphors, to two knowledge engineers experts in conceptual blending and conceptual metaphor theories. We evaluate the representation’s adequacy, correctness, and completeness on a Likert scale from 1 to 7 concerning detected property, source, target and element mappings. The results highlight a consistent pattern where textual components generally outperform their visual counterparts across all evaluated metrics. Specifically, correctness and adequacy in text scored at 5.1 and 6.3, suggesting a positive assessment, whereas the visual aspects scored slightly lower at 4.9 and 5 respectively, indicating room for improvement. The completeness of the text was rated at 5.2, while the visual completeness was lower at 4.4, showing a need for better integration or representation of visual elements to match the textual analysis. Overall, these results suggest a positive performance with a disparity between text and visual outputs, probably because the chosen examples were the ones taken by Petridis et al. [32] and they weren’t further annotated with an RDF syntax to be consistent with the human evaluation.
VII-B1 Error analysis
While metaphor detection shows good performance in known datasets, for conceptual metaphor understanding, errors of the best performing LAG configuration follow the categorization proposed in Wachowiak and Gromann  [40] with the addition of a new category found, the incorrect switching of source and target with respect to the gold standard annotation. Furthermore, generic conceptual metaphors tend to be detected more often than scientific ones.
In particular, misalignment in conceptual metaphor mapping is predominantly due to wrong subelement mapping, which accounts for approximately 56.5% in the WG dataset and 57.1% in the BCTMD. Annotations that are too specific contribute significantly as well, with 23.6% and 28.6% of errors in WG and BCTMD respectively, while overly general descriptions, appearing in 9.3% of WG and 14.3% of BCTMD cases, showing difficulty in identifying the correct taxonomical level of the metaphor. Additionally, less frequent errors, such as switching the source and target (3.24% in WG) and identifying as literal something that was a metaphor (6.94% in WG), show further layers of complexity in the metaphor mapping process, such as misunderstanding the role of source and target domains and determining the metaphoricity level (a sentence might seem literal when the perceived conventionality of the metaphor is high).
These results underscore the need to train and test automatic methods on domain-specific figurative datasets, but also that analogical reasoning is generally not properly performed in these tasks. On the other hand, visual metaphor understanding shows different kinds of errors. The error analysis still follows the one done with human participants in Petridis et al.  [32], with errors involving incorrect objects (21% of the errors). For example, a LLM will identify the shape of a gun as a whistle, or pink Nike shoes as fishes, usually skewing the interpretation. The most common error, however, is incorrect property (57% of the errors): if the LLM correctly identifies the objects, source and target, but still misinterprets the property. For example, in the case of the blending of the car keys and the gun (displayed in Figure LABEL:fig:visualmet), with the annotated property being dangerous, the predicted property is powerful. This error is also observed in human participants and tied to their values and beliefs over the objects displayed in the blend, a further area of investigation. Finally, in 21% of cases, even when the LLM identifies source, target and connecting property, the role mapping explanation can show a misunderstanding of the metaphor (incorrect target symbol). Thus, the evaluation of the property alone as proposed in previous studies [32] is not sufficient to see if the system has understood a metaphor.
The error analysis shows the need to incorporate contextual meaning into the understanding of the metaphor, but also indicates that complete metaphor understanding requires a deeper level of semantic and cultural awareness that despite guidance towards relational analogical thinking, current systems struggle to achieve, also due to the absence of standardized gold-standard annotations that would establish clear metaphorical hierarchies. This multifaceted challenge highlights the gap between computational pattern recognition and the cognitive processes underlying human metaphor comprehension.
VII-C Application of LAG to metaphor generation
In addition to detecting existing metaphors, our framework offers a structured approach to generating new effective metaphors. By formalizing the process of blending conceptual frames, the system can suggest novel metaphors by creatively reconfiguring known concepts into new, plausible configurations according to a specified ontology. This aspect of the framework holds particular promise for applications in creative AI fields, such as advertising, literature, and automated content generation. In future research, we aim to develop a method for generation that addresses limitations found in other studies, such as that described in Chakrabarty et al.  [61]. The authors in this study prompt models to generate visual metaphors using generic, implicit elements, but do not specify how these should be effectively combined to produce meaningful metaphors. Indeed, one reported error in the study was the omission of the target domain or essential elements that were not explicitly requested in the image. Finding the LLM’s errors in understanding visual metaphors can thus improve its ability to generate metaphors in a closer way to our thought processes. This enrichment can foster narrative creation, offering fresh and impactful ways to communicate complex ideas or emotions, and enhancing the engagement and relatability of AI-generated content. Furthermore, with the human-validated data about LLM-generated overarching concepts, scores information can provide some input for the finetuning of weighted blending representations.
VII-D Interpretability of metaphorical datasets
The proposed framework’s explainability through XKG creation helps to outline some aspects that so far haven’t fully emerged in computational approaches to metaphor.
VII-D1 Context and culture
Our research demonstrates that annotating the source and target elements in computational metaphor analysis is insufficient for fully grasping metaphorical meaning. For instance, as illustrated in Figure LABEL:fig:visualmet, the LLM accurately identifies all objects within the image that form the metaphor. However, although the blending property is annotated as “dangerous”, the LLM interprets it as “power”. In the absence of external context that can further frame the reference elements needed for comprehension, both interpretations are arguably valid.
This and other examples from our experiments reveal two key insights. First of all, the perspective adopted by an LLM in interpreting a metaphor reflects its underlying cultural background. In this case, the dual perception of a gun as either powerful or dangerous underscores how cultural context shapes interpretation.
Furthermore, for systems (and users) to reliably converge on metaphorical annotated properties, providing additional context is essential. This requirement for context may also help explain the relatively poor performance observed in human responses, such as distinguishing whether an advertisement is a public campaign, a comic strip, or a standard advertisement.
VII-D2 The dynamics of metaphor annotation
Beyond culture and context-dependence, the human annotation showed that some sentences’ gold standard metaphors in the WG, BCTMD and the Visual Metaphors dataset had multiple possible correct answers or even multiple answers in the same sentence. For example, for the sentence “This mini-controversy erupted when Republicans introduced a string of amendments in a final effort to obstruct passage of the reconciliation bill” the gold-standard metaphor is inhibitting enactment of legislation is impeding motion, while the predicted one is controversy is volcano eruption, which has been noted as correct by the evaluators but would be deemed as incorrect with automatic assessment.
This phenomenon shows the need to keep both the human and the metaphor interpretation process flexible. However, current annotated metaphorical datasets only contain one possible gold standard answer. Future work should address the possibility to incorporate multiple correct answers for a reference text.
VII-E Limitations
Despite its strengths, the current framework has possible improvements. Experiments are limited to English; more text and image data are needed for a more comprehensive assessment. While the work on conceptual metaphors shows some correlation with automated measures, manual evaluation remains essential. Indeed, another survey on the quality of the augmented graphs with a higher quantity of graphs and participants is needed. Additionally, the model’s computational complexity and resource demands may cause time constraints with large datasets or real-time applications. Ongoing refinement is thus necessary to improve scalability and efficiency.
VIII Conclusion and Future Work
In this paper, we presented a Logic-Augmented Generation-based framework that automatically generates extended knowledge graph triples from a base graph using ontology prompting heuristics to enable effective analogical reasoning on various types of unlabeled figurative data. We gathered four datasets for metaphor detection and understanding and introduced a novel dataset featuring domain-specific scientific conceptual metaphors. Our results in multimodal metaphor detection and understanding outperform current baselines with Claude Sonnet 3.5 and even surpass human performance in comprehending visual metaphors, although the accuracy for scientific metaphors remains significantly lower than for generic ones. The error analysis confirms recent claims that, while LLMs excel at surface associations, they still lack robust relational reasoning processes, partly due to insufficient contextual explanations in existing datasets. Despite this shortcoming, which calls for more diverse and fine-grained metaphorical data, leveraging the Blending Ontology to generate augmented knowledge graphs enhances output explainability compared to current methods, which often fail to clarify why, even when the objects of a metaphor are detected, the metaphorical property is not correctly identified. The proposed approach therefore advances computational models for multimodal figurative language understanding in unstructured data, with significant implications for AI in natural language processing, creativity, and applications such as hate speech detection. Future work will focus on refining the framework’s computational complexity and testing it on a broader range of annotated data.